# Chapter 0 Hadron Production Processes

### v1

@@ 2 Perturbative Treatments: Photon-Hadron Channels and $K$-matrix Born Approximation
The much weaker electromagnetic coupling constant, $e^{2}\sim\alpha_{f}\simeq 1/137\ll g^{2}_{\pi N}\sim 14$, the photo-production reaction channels may be treated perturbatively in leading order of the $\gamma N$, $\gamma N^{\ast}$, and photon-hadron vertices in general. Thus, the T-matrix elements involving a photon are treated in the lowest order Born approximation, $T_{\gamma h}\to V_{\gamma h}$, where $h$ denotes a hadron.
A significant reduction of the numerical effort is achieved by the so-called $K$-matrix Born approximation by using $K\equiv V$ [90]. This approximation corresponds to neglecting the real part of $G_{BS}$. Hence, it is in fact a pole approximation since the intermediate propagation is frozen to the on-shell contributions produced by the remaining imaginary part of $G_{BS}$.
The systematic term-by-term summation of scattering series and similar perturbative series goes back to Max Born (1882-1970). Max Born shared the 1954 Nobel Prize in Physics with Walther Bothe (1891-1957) for his fundamental research in quantum mechanics, especially in the statistical interpretation of the wave function. Moreover, Max Born’s book of 1933 on Optik [91] played a central role in theory and experiments on gamma spectroscopy. In 1959, the book was updated, extended by contributions from other authors, and republished in the English language, experiencing several reprints [92].
@@ 3 Interaction Potential and Scattering Matrix
The hadron-hadron interaction matrix $V$ is built by a sum of $s$-, $u$-, and $t$-channel Feynman diagrams (once invented by Richard Feynman for visualizing QFT/QED amplitudes) of the type shown in Fig. 11. They are derived from the underlying effective model Lagrangians, e.g., see [93] for the one used in GiM. In order to reduce the number of model parameters, the non-resonant background terms should be derived consistently by the $u$- and $t$-channel diagrams resulting from the model Lagrangian.
Figure 11: Structure of the tree-level interaction potential $V$. $s-$, $u-$, and $t-$ channel interactions defining the non-resonant background contributions are shown in the first line, including contact terms which are chosen such that gauge invariance is assured. The so-called z-diagrams, displayed in the second line are generic for the double-pion channels. $s-$ channel resonance interactions are depicted in the last line. Time is running from left to right.
In a symbolical short-hand notation, the bare, Born-term interaction for a reaction $a\to b$ at center-of-mass energy $w=\sqrt{s}$ is written as
@@
where the contributions represent the first, the second, and the third line of Fig.11, respectively, where
@@
In each partial wave channel $[JLP]$ a hadronic interaction matrix is found
@@
where $V_{\pi\pi}\equiv V_{\pi B\pi B^{\prime}}$ for baryons $B,B^{\prime}=N,N^{*},\Lambda,\Sigma\ldots$, and correspondingly for the other meson channels..
@@
where also the K-matrix Born approximation is indicated. The phase space factor $\rho_{C}$ is produced by evaluating the integral over the delta-distributions of $Im[G_{BS}]$.
@@
4 Couple Channels Projects for Partial-Wave Analyses
@@
Historically, Adrien-Marie Legendre (1752-1833) was the first to find the multipole expansion of the gravitational two-body potential [94] for which he had to invent his famous system of orthogonal polynomials, since then indispensable for any partial wave expansion. Rayleigh’s book on the Theory of Sound [95] established PWA as a tool in classical wave mechanics. In quantum scattering theory PWA is the mathematical tool to calculate reaction amplitudes for given interactions. On the practitioner’s side PWA is the key technique for determining reaction amplitudes by fits to scattering data and concluding on the underlying interaction. That task is in fact a non-trivial mathematical problem, corresponding to the solution of an in general ill-posed inversion problem - as was pointed out some time ago on quantum-mechanical grounds, e.g., by Andrei Nikolaevich Tikhonov [96] (1906-1993).
@@
Over the years, in hadron physics several CC approaches have been developed, out of which a few long-term projects have emerged. The investigations of Höhler’s [97] and Cutkosky’s group [98] have the merits of establishing computational methods in hadron spectroscopy as tools for systematically enlarging and interpreting the database. Coupled-channel (CC) approaches have been proven to be an efficient tool to extract hadron properties from experiments. They are firmly established as indispensable workhorses for research on hadron spectroscopy, playing the dominant role, especially in baryon spectroscopy.
@@
In that tradition, the SAID project is probably the one with the largest long-term impact on the community. SAID, initiated by Richard A. Arndt, L. David Roper, and their group at Virginia Polytechnic Institute and State University and since then fostered continuously at The George Washington University since 1999 [99], provides compilations of experimental data, constant updates of CC methods and results. The easy online accessibility is an important service to the community and strongly supports the research on hadron spectroscopy. Unlike the other CC approaches referred to below, SAID does not assume ad-hoc resonance contributions but derives partial wave amplitudes directly from data, followed by an analysis on pole positions in the complex energy plane and BW parameters. The latter approach has natural limits because of: (i) revealing wide resonances up to a width $\Gamma<500~{}\mathrm{MeV}$, (ii) missing transitions of small Branching Ratios $BR<4\%$, and (iii) tending (by construction) to miss narrow resonances if $\Gamma<20~{}\mathrm{MeV}$. However, in SAID narrow resonances are accounted for by a modified PWA method [100, 101]. The computational SAID approach accounts for energy dependencies of self-energies on the basis of the Chew-Mandelstam method.
@@
The MAID project, connected to the MAMI facility at Johannes Gutenberg University of Mainz, has been following similar routes. The Bonn-Gatchina (BnGa) PWA project of Bonn University and Saint Petersburg Nuclear Physics Institute (PNPI) at Gatchina [^17] Gatchina is a town in $40~{}\mathrm{km}$ on the south of Saint Petersburg and by about $700~{}\mathrm{km}$ far from Moscow. collaboration and centered at the ELSA laboratory at Bonn University is another influential CC activity [102]. BnGa utilizes the so-called N/D approach (see [103]) and otherwise incorporates loop corrections into the K-matrix approach in a spirit similar to SAID. The BnGa, MAID, and the SAID solutions are accessible online.
@@
The Jülich-Bonn-Washington Collaboration (Jü-Bo), including Jülich Center für Hadron Physics, Bonn University, and The George Washington University, has been formulating a coupled channels model [104, 105]. That approach, in fact, is based largely on the work of Haberzettl et al. [106].
@@
Over a long period, the Giessen group at Justus Liebig-University of Giessen has been developing its own covariant CC model, the Giessen Model (GiM). Initiated in the early 1990s by Ulrich Mosel, the project has grown over a few decades to a full-fledged and versatile tool for baryon spectroscopy by meson photoproduction on the nucleon and subsequent spectroscopy of the decay channels. The GiM utilizes a K-matrix approach strictly oriented on Lagrangian methods with special attention on the consistency of interactions [107]. The theoretical background of GiM is closely oriented to the discussions in the previous paragraphs. Applications range from single pion, eta, and kaon production and investigations of vector meson production to the population of two-pion channels and Compton scattering [108] [^18] Named after Arthur Holly Compton (1892-1962) who won the 1927 Nobel Prize in Physics for his discovery of the Compton effect, which demonstrated the particle nature of electromagnetic radiation.. A recent overview of the GiM is found in [109].
@@ 5 Quantum Interference in Hadron Spectra
At the time of writing this article, the science and technology world is celebrating 100 years of quantum mechanics. Among others, coherence of wave functions and interference of matrix elements are defining properties of quantum systems. While in atomic, molecular, and solid state systems quantum interference is used actively to manipulate the systems, see, e.g.[110], the vastly different conditions encountered in nuclear and sub-nuclear systems usually inhibit direct actions at a comparable level. Important signatures for interference phenomena are irregularities in line shapes, observed as significant deviations from Lorentz (named after Hendrik Antoon Lorentz (1853-1928) [^19] Lorentz shared the 1902 Nobel Prize in Physics with Pieter Zeeman (1865-1943) for their discovery and theoretical explanation of the Zeeman effect (named after Pieter Zeeman).) (or Gaussian (named after Carl Friedrich Gauss (1777-1855)) shapes as expected for isolated, non-interacting resonances. In nuclear spectra, exceptional line shapes have been seen in a few cases [111, 112, 113] and interpreted in the Fano formalism [^20] Ugo Fano (1912-2001), originally a member of Fermi’s team, was a master at understanding how radiation interacts with matter. His work set the agenda for much of modern atomic physics, and had a broad sweep across field as stated in an obituary published in [114]. [115], being used also in atomic and laser physics in [110]. In [116, 117] that formalism was applied to interpret irregular line shapes of the $D\bar{D}$ decay modes of the aforementioned $\psi(3770)$ resonance.
@@
Though the phenomenon of quantum-mechanical interference has been known for many years, there are still open questions, not only with respect to hadron physics but also in electronic systems, see, e.g.[118]. In the review [119], Azimov discussed how the interference of resonances may and does work. A rich source is data on rare decay modes of well-known resonances which demonstrate a wide variety of possible different manifestations of interference. Some special kinds of resonance interference, not yet sufficiently studied and understood, are also briefly considered. The interference may give useful experimental procedures to search for new resonances with arbitrary quantum numbers, even with exotic ones, and to investigate their properties.
@@
In Figure 12 resonance interference is illustrated for $\rho^{0}$, $\omega$-, and $\phi$-meson production in the reaction $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}$. Total cross sections collected by the SND Collaboration are shown. Actually, for the $3\pi$ decay case, $\Gamma(\rho^{0}\to 3\pi)=0.015~{}\mathrm{MeV}$ (Isospin symmetry violation), $\Gamma(\omega\to 3\pi)=7.58~{}\mathrm{MeV}$, and $\Gamma(\phi\to 3\pi)=0.65~{}\mathrm{MeV}$ (Zweig rule violation) [38]. It has a clear BW-like peak in the ($\rho^{0}$, $\omega$) region and a bump–dip structure in the $\phi$ region. Background near $\phi$ changes slowly vs. nearly standard interference curve. Instead of $\phi$–peak, there are both bump and dip, each has a form different from BW; max/min shifts from $\phi$-mass. There is a similar shift for the $\rho$-mass. Then $\rho$- contribution here deforms the $\omega$-tail.
@@
The interpretation of the recent BESIII cross sections for the reaction $e^{+}e^{-}\to J/\psi\pi^{+}\pi^{-}$ [121] as an observation of the $X(3872)$ resonance [122] via a destructive interference between small resonance and large background contributions (that is a natural way for the inelastic reaction, see, for instance, [123]) (Fig. 13). The large non-resonance contribution magnifies a small resonance effect, as described in different contexts by the Fano mechanism. The observed dip is small, but visible, and more statistics would be highly desirable.
@@
The data displayed in Figs. 12 and 13 demonstrate the direct interference between two resonances decaying into the same final states [^21] A while ago Richard Feynman said: When looking at Maxwell equations, it is hard to imagine how beautiful the rainbow is. Then Yakov Azimov (1938-2016) extended it: Everybody knows that the interference does exist. But it is not always easy to imagine how it will work in a particular case..
8 Meson Photoproduction
@@
The determination of the resonance properties for all accessible baryon states is a central objective in nuclear physics. The extracted resonance parameters provide a crucial body of data for understanding the dynamics of nucleon excitations and the spectral distributions, essential for testing o´phenomenological models of the nucleon, calibrating lattice QCD calculations as a way of connecting non-perturbative hadron physics to QCD. The spectrum of $N^{\ast}$ and $\Delta^{\ast}$ non-strange baryon resonances of masses up to about 2 GeV is probably the best-studied sector of hadron physics. Meson-nucleon scattering and photoproduction of mesons on nucleon and nuclear targets have led to a wealth of data, lately also supplemented by combined systematic studies of kaon and associated hyperon production experiments. Properties of the known resonances continue to become better determined as experiments involving polarized beams, targets, and recoil measurements are expanded and refined. With increasing energy, multi-meson channels will open. They provide the opportunities to identify states coupled weakly to two-body channels like $\pi+N$. Since production reactions involving the octet - neglecting the $\eta^{\prime}$ singlet - of pseudo-scalar mesons and the related baryon resonances are studied the best, they will be in the center of the following discussions. To a lesser degree, also production reactions leading to the nonet of vector mesons have been investigated and will be addressed selectively. Thus, in the following sections, reactions involving the baryon octet and the decuplet, Fig. 6 and two of the three meson nonets of Fig. 7 will be discussed.
The scalar mesons are suffering from short lifetimes and ill-defined spectral distributions, see, e.g., the status of the $f_{0}(500)/\sigma$ meson [38]. Moreover, quark model studies predict that the scalar mesons of masses below $1~{}\mathrm{GeV}$ are most likely tetra-quark states of $qq-\bar{q}\bar{q}$ structure [124]. Therefore, light scalar meson production on the nucleon will not be considered. Accordingly, reactions with direct multi-meson production will be neglected, too.
1 Single Pseudoscalar Meson Photoproduction
Pion Photoproduction on Proton

### v2

@@ 2 Perturbative Treatments: Photon-Hadron Channels and $K$-matrix Born Approximation
The much weaker electromagnetic coupling constant, $e^{2}\sim\alpha_{f}\simeq 1/137\ll g^{2}_{\pi N}\sim 14$, allows the photo-production reaction channels to be treated perturbatively in the leading order of the $\gamma N$, $\gamma N^{\ast}$, and photon-hadron vertices in general. Thus, the T-matrix elements involving a photon are treated in the lowest order Born approximation, $T_{\gamma h}\to V_{\gamma h}$, where $h$ denotes a hadron.
A significant reduction of the numerical effort is achieved by the so-called $K$-matrix Born approximation by using $K\equiv V$ [103]. This approximation corresponds to neglecting the real part of $G_{BS}$. Hence, it is, in fact, a pole approximation since the intermediate propagation is frozen to the on-shell contributions produced by the remaining imaginary part of $G_{BS}$.
The systematic term-by-term summation of scattering series and similar perturbative series goes back to Max Born (1882-1970). Max Born shared the 1954 Nobel Prize in Physics with Walther Bothe (1891-1957) for his fundamental research in quantum mechanics, especially in the statistical interpretation of the wave function. Moreover, Max Born’s book of 1933 on Optik [104] played a central role in the theory and experiments on gamma spectroscopy. In 1959, the book was updated, extended by contributions from other authors, and republished in English, experiencing several reprints [105].
@@ 3 Interaction Potential and Scattering Matrix
The hadron-hadron interaction matrix $V$ is built from a sum of $s$-, $u$-, and $t$-channel Feynman diagrams (originally introduced by Richard Feynman for visualizing QFT/QED amplitudes) of the type shown in Fig. 11. They are derived from the underlying effective model Lagrangians, e.g., see [106] for the one used in the Giessen Model (GiM). In order to reduce the number of model parameters, the non-resonant background terms should be derived consistently from the $u$- and $t$-channel diagrams resulting from the model Lagrangian.
Figure 11: Structure of the tree-level interaction potential $V$. $s$-, $u$-, and $t$- channel interactions defining the non-resonant background contributions are shown in the first line, including contact terms which are chosen such that gauge invariance is assured. The so-called z-diagrams, displayed in the second line, are generic for the double-pion channels. $s$- channel resonance interactions are depicted in the last line. Time is running from left to right.
In symbolic short-hand notation, the bare Born-term interaction for a reaction $a\to b$ at center-of-mass energy $w=\sqrt{s}$ is written as
@@
where the contributions represent the first, second, and third lines of Fig. 11, respectively, where
@@
In each partial wave channel $[JLP]$, a hadronic interaction matrix is found
@@
where $V_{\pi\pi}\equiv V_{\pi B\pi B^{\ast}}$ for baryons $B,B^{\ast}=N,N^{\ast},\Lambda,\Sigma\ldots$, and correspondingly for the other meson channels..
@@
where the K-matrix Born approximation is also indicated. The phase space factor $\rho_{C}$ is produced by evaluating the integral over the delta-distributions of $\mathrm{Im}[G_{BS}]$.
@@
4 Coupled Channel Projects for Partial-Wave Analyses
@@
Historically, Adrien-Marie Legendre (1752-1833) was the first to find the multipole expansion of the gravitational two-body potential [107] for which he had to invent his famous system of orthogonal polynomials, since then indispensable for any partial wave expansion. Rayleigh’s book on the Theory of Sound [108] established PWA as a tool in classical wave mechanics. In quantum scattering theory, Partial Wave Analysis (PWA) is the mathematical tool used to calculate reaction amplitudes for given interactions. On the practitioner’s side, PWA is the key technique for determining reaction amplitudes by fitting scattering data and concluding on the underlying interaction. That task is, in fact, a non-trivial mathematical problem corresponding to the solution of, in general, an ill-posed inversion. problem, as was pointed out some time ago on quantum-mechanical grounds, e.g., by Andrei Nikolaevich Tikhonov [109] (1906-1993).
@@
Over the years, in hadron physics, several CC approaches have been developed, out of which a few long-term projects have emerged. The investigations of Höhler’s [110] and Cutkosky’s group [111] have the merit of establishing computational methods in hadron spectroscopy as tools for systematically enlarging and interpreting the database. Coupled-channel (CC) approaches have proven to be an efficient tool for extracting hadron properties from experiments. They are firmly established as indispensable workhorses for research on hadron spectroscopy, playing a dominant role, especially in baryon spectroscopy.
@@
In that tradition, the SAID project is probably the one with the largest long-term impact on the community. SAID, initiated by Richard A. Arndt, L. David Roper and their group at Virginia Polytechnic Institute and State University has since continuously fostered research at The George Washington University since 1999 [112], providing compilations of experimental data, constant updates of CC methods, and results. The easy online accessibility is an important service to the community and strongly supports research on hadron spectroscopy. Unlike the other CC approaches referred to below, SAID does not assume ad-hoc resonance contributions but derives partial wave amplitudes directly from data, followed by an analysis of pole positions in the complex energy plane and Breit-Wigner parameters. The latter approach has natural limits because of: (i) revealing wide resonances up to a width $\Gamma<500~\mathrm{MeV}$, (ii) missing transitions of small Branching Ratios $BR<4\%$, and (iii) tending (by construction) to miss narrow resonances if $\Gamma<20~\mathrm{MeV}$. However, in SAID, narrow resonances are accounted for by a modified PWA method [113, 114]. The computational SAID approach accounts for energy dependencies of self-energies based on the Chew-Mandelstam method.
@@
The MAID project, connected to the MAMI facility at Johannes Gutenberg University of Mainz, has been following similar routes. The Bonn-Gatchina (BnGa) PWA project of Bonn University and the Saint Petersburg Nuclear Physics Institute (PNPI) at Gatchina [^18] Gatchina is a town in $40~\mathrm{km}$ south of Saint Petersburg and by about $700~\mathrm{km}$ far from Moscow. Collaboration and centered at the ELSA laboratory at Bonn University is another influential CC activity [115]. BnGa utilizes the so-called N/D approach (see [116]) and otherwise incorporates loop corrections into the K-matrix approach in a spirit similar to SAID. The BnGa, MAID, and SAID solutions are accessible online.
@@
The Jülich-Bonn-Washington Collaboration (Jü-Bo), including Jülich Research Center, Bonn University, and The George Washington University have been formulating a coupled channels model [117, 118]. That approach, in fact, is based largely on the work of Haberzettl et al. [119].
@@
Over a long period, the Giessen group at Justus Liebig-University of Giessen has been developing its own covariant CC model, GiM. Initiated in the early 1990s by Ulrich Mosel, the project has grown over a few decades into a full-fledged and versatile tool for baryon spectroscopy through meson photoproduction on the nucleon and subsequent spectroscopy of the decay channels. The GiM utilizes a K-matrix approach strictly oriented toward Lagrangian methods, with special attention to the consistency of interactions [120]. The theoretical background of GiM is closely oriented to the discussions in the previous paragraphs. Applications range from single pion, eta, and kaon production and investigations of vector meson production to the population of two-pion channels and Compton scattering [121] [^19] Named after Arthur Holly Compton (1892-1962), who won the 1927 Nobel Prize in Physics for his discovery of the Compton effect, which demonstrated the particle nature of electromagnetic radiation.. A recent overview of the GiM is found in [83].
@@ 5 Quantum Interference in Hadron Spectra
At the time of writing this article, the science and technology world is celebrating 100 years of quantum mechanics. Among others, coherence of wave functions and the interference of matrix elements are defining properties of quantum systems. While in atomic, molecular, and solid state systems, quantum interference is actively used to manipulate the systems (see, e.g., Ref. [122]), the vastly different conditions encountered in nuclear and sub-nuclear systems usually inhibit direct actions at a comparable level. Important signatures for interference phenomena are irregularities in line shapes, observed as significant deviations from Lorentz (named after Hendrik Antoon Lorentz (1853-1928) [^20] Lorentz shared the 1902 Nobel Prize in Physics with Pieter Zeeman (1865-1943) for their discovery and theoretical explanation of the Zeeman effect (named after Pieter Zeeman).) (or Gaussian (named after Carl Friedrich Gauss (1777-1855)) shapes, as expected for isolated, non-interacting resonances. In nuclear spectra, exceptional line shapes have been seen in a few. cases [123, 124, 125] and interpreted in the Fano formalism [^21] Ugo Fano (1912-2001), originally a member of Fermi’s team, was a master at understanding how radiation interacts with matter. His work set the agenda for much of modern atomic physics, and had a broad sweep across the field, as stated in an obituary published in [126]. [127], being used also in atomic and laser physics in [122]. In [128, 129], that formalism was applied to interpret irregular line shapes of the $D\bar{D}$ decay modes of those as mentioned earlier $\psi(3770)$ resonance.
@@
Though the phenomenon of quantum-mechanical interference has been known for many years, there are still open questions, not only with respect to hadron physics but also in electronic systems, see, e.g., Ref. [130]. In the review [131], Azimov discusses how the interference of resonances may and does work. A rich source of data is on rare decay modes of well-known resonances, which demonstrate a wide variety of possible different manifestations of interference. Some special kinds of resonance interference, which have not yet been sufficiently studied and understood, are also briefly considered. The interference may provide useful experimental procedures to search for new resonances with arbitrary quantum numbers, including exotic ones, and to investigate their properties.
@@
In Figure 12 resonance interference is illustrated for $\rho^{0}$, $\omega$-, and $\phi$-meson production in the reaction $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}$. Total cross sections collected by the SND Collaboration are shown. For the $3\pi$ decay case, $\Gamma(\rho^{0}\to 3\pi)=0.015~\mathrm{MeV}$ (Isospin symmetry violation), $\Gamma(\omega\to 3\pi)=7.58~\mathrm{MeV}$, and $\Gamma(\phi\to 3\pi)=0.65~\mathrm{MeV}$ (Zweig rule violation) [41]. It has a clear BW-like peak in the ($\rho^{0}$, $\omega$) region and a bump–dip structure in the $\phi$ region. The background near $\phi$ changes slowly compared to the nearly standard interference curve. Instead of the $\phi$–peak, there are both bumps and dips, each having a form different from BW; max/min shifts from $\phi$-mass. There is a similar shift for the $\rho$-mass. Then, the $\rho$ contribution here deforms the $\omega$-tail.
@@
The recent BESIII cross section data for the reaction $e^{+}e^{-}\to J/\psi\pi^{+}\pi^{-}$ [133] have been interpreted as an observation of the $X(3872)$ resonance [134] via a destructive interference of a sharp resonance and large background contributions (this is a natural way for the inelastic reaction, see, for instance, [135]) (Fig. 13). The large non-resonance contribution magnifies a small resonance effect, as described in different contexts by the Fano mechanism [125]. The observed dip is small, but visible, and more statistics would be highly desirable.
@@
The data displayed in Figs. 12 and 13 demonstrate the direct interference between two resonances decaying into the same final states [^22] A while ago, Richard Feynman said: When looking at Maxwell equations, it is hard to imagine how beautiful the rainbow is.. Then Yakov Azimov (1938-2016) extended it: Everybody knows that the interference does exist. But it is not always easy to imagine how it will work in a particular case..
8 Single Meson Photoproduction
1 Prelude
@@
The determination of the resonance properties for all accessible baryon states is a central objective of hadron and nuclear physics. The extracted resonance parameters provide a crucial body of data for understanding the dynamics of nucleon excitations and the spectral distributions, essential for testing phenomenological models of the nucleon, calibrating lattice QCD calculations as a way of connecting non-perturbative hadron physics to QCD. The spectrum of $N^{\ast}$ and $\Delta^{\ast}$ non-strange baryon resonances with masses up to about $2~\mathrm{GeV}$ is probably the best-studied sector of hadron physics. Meson-nucleon scattering and photoproduction of mesons on nucleon and nuclear targets have led to a wealth of data, which has also been supplemented lately by combined systematic studies of kaon and associated hyperon production experiments. Properties of the known resonances continue to become better determined as experiments involving polarized beams, targets, and recoil measurements are expanded and refined. With increasing energy, multi-meson channels will open. They provide the opportunity to identify states that are weakly coupled to two-body channels, such as $\pi+N$. Since production reactions involving the octet-neglecting the $\eta^{\prime}$ singlet of pseudo-scalar mesons and the related baryon resonances are studied the best, they will be the focus of the following discussions. To a lesser degree, production reactions leading to the nonet of vector mesons have also been investigated and will be addressed selectively. Thus, in the following sections, reactions involving the baryon octet and the decuplet, Fig. 6 and two of the three meson nonets of Fig. 7 will be discussed.
Experimentally, light scalar mesons typically suffer from short lifetimes and ill-defined spectral distributions; see,e.g., the experimental status of the $f_{0}(500)/\sigma$ meson [41]. However, after decades of research, the properties of the lowest scalar meson are encapsulated as discussed comprehensively in [136] and reviewed in detail in a separate article of the encyclopedia [137]. Diffuse objects such as $f_{0}(500)$ remain a challenge for coupled channel approaches because, in principle, they require solving at least a numerically very involved coupled three-body problem in order to properly account for the cross-talk between the various interacting many-particle channels. A brief account of sigma meson production on the nucleon is given in one of the following sections in the context of double pion production.
2 Single Pion Photoproduction
Pion Photoproduction on the Proton